OpenStudy (anonymous):

Question 3 : Signals : How this is true? I am not getting it properly.. \[\large \color{green}{u[n] = \sum_{k \rightarrow - \infty}^{n} \delta[k]}\]

OpenStudy (anonymous):

\(u[n]\) is Unit Step Discrete signal \(\delta[k]\) is Unit Impulse or Sample Sequence..

OpenStudy (anonymous):

I think I have got this.. :)

OpenStudy (anonymous):

First case: For n < 0.. 

OpenStudy (anonymous):

Second Case: \(n \ge 0\) 

OpenStudy (anonymous):

As : It is equivalent to : \( u[n] = \begin{cases} 1, && n \ge 0 \\ 0, && n <0\end{cases}\) So: \[\large \color{blue}{\sum_{k \rightarrow - \infty}^{n} \delta[k] = u[n]}\]